Find the ratio of x to y: 4/5=x-y/x+y
Answer by stanbon(75887)

Find the ratio of x to y: 4/5=x-y/x+y
------
4(x+y) = 5(x-y)
----
4x + 4y = 5x-5y
----
x = 9y
---
x/y = 1/9
-------------
Cheers,
Stan H.
--------------